OpenStudy (anonymous):

somebody please help solve : 3x/(2) + 1/(x)= 3x/(4)

OpenStudy (anonymous):

what does the question ask to do? solve for x?

OpenStudy (anonymous):

it just says solve so i would guess for x

OpenStudy (anonymous):

im not getting any solutions...

OpenStudy (anonymous):

1/x = 3x/4 - 6x/4 1/x = -3x/4 4 = -3x² 0 = -3x² - 4 this function never equals 0 = no value for x...

OpenStudy (anonymous):

is it possible for the answer to equal i radical 4/3?

OpenStudy (anonymous):

yea unless u can have i in the answer lol

OpenStudy (anonymous):

i guess we can

OpenStudy (anonymous):

can you look at this simplify the expression: 5 +\[\sqrt{7}\]\[\div\]5-\[\sqrt{7}\]

OpenStudy (anonymous):

lol wtf? that came out all weird.....repost?

OpenStudy (anonymous):

it should read 5+ radical 7 divided by 5- radical 7

OpenStudy (anonymous):

and it needs to be simplified